# EMLoC: Emulator-based Memory-efficient Fine-tuning with LoRA Correction

### v1

@@ Abstract
Open-source foundation models have seen rapid adoption and development, enabling powerful general-purpose capabilities across diverse domains. However, fine-tuning large foundation models for domain-specific or personalized tasks remains prohibitively expensive for most users due to the significant memory overhead beyond that of inference. We introduce EMLoC, an Emulator-based Memory-efficient fine-tuning framework with LoRA Correction, which enables model fine-tuning within the same memory budget required for inference. EMLoC constructs a task-specific light-weight emulator using activation-aware singular value decomposition (SVD) on a small downstream calibration set. Fine-tuning then is performed on this lightweight emulator via LoRA. To tackle the misalignment between the original model and the compressed emulator, we propose a novel compensation algorithm to correct the fine-tuned LoRA module, which thus can be merged into the original model for inference. EMLoC supports flexible compression ratios and standard training pipelines, making it adaptable to a wide range of applications. Extensive experiments demonstrate that EMLoC outperforms other baselines across multiple datasets and modalities. Moreover, without quantization, EMLoC enables fine-tuning of a 38B model on a single 24GB consumer GPU—bringing efficient and practical model adaptation to individual users.
@@ 1 Introduction
General-purpose foundation models have demonstrated impressive zero-shot capabilities across a wide range of benchmarks [11, 8, 2, 35, 6]. For real-world deployment, such as domain-specific tasks [21, 42] or personalized user behavior [26, 30] , further customization by fine-tuning is still required. However, fine-tuning typically incurs significantly more memory overhead than inference [28]. Consequently, if users have a fixed amount of available computing resources, they will be forced to choose between two unfavorable options. First, they can use a small model that fits within their memory budget for fine-tuning as shown in Fig. 1(a), but this sacrifices the emergent capabilities [38] of larger models and underutilizes hardware during inference. Alternatively, they can opt for a large model that fully utilizes resources during inference but exceeds memory limits for fine-tuning as shown in Fig. 1(b), making user-specific adaptation infeasible and potentially limiting performance in specialized applications. This paper addresses a central research question: Is it possible to design a fine-tuning strategy such that users can fine-tune a model under the same memory budget as inference?
@@
The memory cost of fine-tuning can be broadly attributed to three components: optimizer states, intermediate activations, and the model parameters themselves, as marked in Fig. 1 with different colors. Initial efforts to reduce the memory usage of fine-tuning concentrated on the first two components. The first component, optimizer states, stores auxiliary information such as momentum and variance in the Adam optimizer [15] for each trainable parameter. This overhead can be mitigated using parameter-efficient fine-tuning (PEFT) methods [13, 20, 41], such as LoRA [13], which reduces trainable parameters by only updating lightweight low-rank matrices. The second component, intermediate activations, refers to the values retained during the forward pass for use in backpropagation. Previous works reduced it by modifying model architectures [34, 10, 19] or gradient checkpointing [4]. However, since the components reduced by both strategies still presented during fine-tuning, reducing them alone, without addressing the memory usage of model parameters, cannot fully close the memory gap between fine-tuning and inference (Fig. 1(b)).
@@
Few recent works have taken initial steps toward addressing the third component, the memory cost of model parameters. For example, Offsite-Tuning [40] proposes to drop some of the intermediate layers entirely during fine-tuning and to update only the top and bottom layers, thereby saving some memory occupied by model parameters. LORAM [45] introduces row-pruning as a method to reduce model parameters during fine-tuning, allowing the updating of the unpruned rows in the intermediate layers. Despite their promise, these methods typically rely on memory-intensive full-model training. This entails a strong and often impractical assumption that well-resourced institutions must be involved, making them unfeasible for individual users. In addition, they impose strict constraints on which subsets of weights can be updated, limiting their flexibility and general applicability across tasks and architectures. Moreover, they overlook the fact that different base models are used during fine-tuning and inference, introducing potential mismatches that degrade performance.
@@
Figure 1: The dilemma caused by additional memory overhead during fine-tuning. (a) Users opt for a smaller 8B model, sacrificing emergent capabilities and underutilizing available hardware. (b) Use of a larger 26B model requiring memory exceeding the hardware limit even with LoRA [13] and gradient checkpointing [4] techniques. (c) Our EMLoC utilizes a smaller model during fine-tuning, allowing the same budget for both training and inference.
@@
In this paper, we propose EMLoC, an Emulator-based Memory-efficient fine-tuning framework with LoRA Correction. EMLoC completely eliminates the memory cost gap between inference and training and can be carried out solely by individual users. The central idea is to perform fine-tuning on a low-rank model which we call the emulator so that the memory taken up by model parameters can be reduced as depicted in Fig. 1(c). To make this approach effective, we introduce an efficient, downstream-aware emulator construction procedure based on activation-aware singular value decomposition (SVD), using a small subset of downstream data for calibration. This yields a lightweight emulator tailored to the downstream data, without requiring costly full-model training. Fine-tuning is then conducted entirely on the emulator without any assumptions about the fine-tuning process—any standard training pipeline can be used. However, since the LoRA modules are fine-tuned on the emulator rather than the original model, naively merging the LoRA modules into the original model will suffer from the misalignment between the full model and the compressed emulator. To mitigate this issue, we further introduce a novel LoRA correction algorithm, which adjusts the parameters to compensate such discrepancy during inference.
@@ 2.1 Reducing Optimizer State
There are mainly two ways to reduce the memory required for optimizer states: by reducing the number of trainable parameters [13, 41, 43, 46, 14, 12, 20], or by reducing auxiliary information required per parameter [31, 18, 49]. A prominent example of the first approach is LoRA [13], which introduces trainable low-rank adapters into frozen pre-trained models to achieve parameter-efficient fine-tuning. For the second approach, Adafactor [31] eliminates the need to store full-rank second-moment estimators by using a factored approximation. While these methods are effective, a non-trivial number of trainable parameters are required to maintain acceptable performance. As a result, the optimizer states associated with those parameters continue to contribute a non-negligible portion of the overall memory footprint.
@@ 2.2 Reducing Intermediate Activation
Another way to fine-tune with less memory overhead is to reduce memory from intermediate activations. Two explored strategies are shortening the backpropagation path [34, 25, 41, 44] and recomputing activations during the backward pass [4, 10, 19]. An example of the first strategy is Ladder Side-Tuning [34], where trainable modules are placed outside the backbone to avoid backpropagation through the backbone. [19] carefully design the position and initialization of the adapter modules so that the pre-trained model becomes reversible, which enables recomputing intermediate activation from the output during backpropagation. However, these methods cannot completely eliminate activation memory due to the fundamental need for gradients, and they typically incur additional computational overhead.
@@ 2.3 Reducing Model Parameters
Beyond optimizing optimizer states and intermediate activations, very recently reducing the memory footprint of the model parameters during fine-tuning has been explored. It is worth noting that quantization-based methods [9, 48] are largely orthogonal to our focus: they do not reduce fine-tuning memory relative to inference, and can therefore be combined with our method. An approach to reducing parameter memory during fine-tuning involves using a pruned model for training while maintaining a full model for inference [40, 45]. For example, LORAM [45] assumes that the model publisher provides a row-pruned and continually pre-trained variant of the original model. Users fine-tune this reduced model and then transfer the learned LoRA modules back to the full model. However, their setting requires continual pre-training of the pruned model, making it inaccessible to individual users. Furthermore, only the unpruned weights can be fine-tuned, limiting flexibility in selecting which parameters to adapt. In contrast, our work introduces a pipeline that can be performed by users individually and allows flexible selection of tunable parameters.
@@ 3 Method
Our proposed method, EMLoC, is a memory-efficient fine-tuning framework that enables individual users to fine-tune a model with the same memory cost as inference. An overview of EMLoC is shown in Fig. 2. EMLoC consists of three main stages. First, we construct a downstream-aware lightweight emulator by performing activation-aware singular value decomposition (SVD) [37] on a small calibration set drawn from the downstream data. Second, we fine-tune the emulator with LoRA modules using any standard training pipeline. The reduced parameter size of the emulator is crucial for enabling memory-efficient fine-tuning. Finally, to address the misalignment issue between the original model and the emulator, we present a novel LoRA correction algorithm to correct the learned LoRA modules, and transfer them back to the original model before inference.
@@ 3.2 Downstream-aware Emulator Construction and Fine-tuning
Specifically, we construct the set of emulator weights $\mathbb{W}^{\mathcal{E}}$ by applying the activation-aware SVD method proposed in SVD-LLM [37] to each weight matrix $W\in\mathbb{W}$
@@
Owing to the specific design choices in the emulator, we can apply any standard PEFT techniques [13, 41, 43] to fine-tune the significantly smaller emulator, achieving memory-efficient fine-tuning. In particular, if we aim to fine-tune a weight matrix $W_{i}$ in the original model, we instead insert a LoRA module next to $W^{\mathcal{E}}_{i}$ to fine-tune the emulator as in conventional frameworks. This allows us to use any existing training pipeline without modification, making our method a plug-and-play replacement for conventional LoRA fine-tuning, while further reducing memory usage.
@@ 5. $W^{c}_{B}\leftarrow W^{\prime}_{B}-\text{clamp}(\Delta,\lambda)$
To fully utilize the capacity of the original model during inference, we aim to transfer the LoRA modules $\Lambda$ fine-tuned on $\mathcal{E}$ back to the original model $\mathcal{M}$. However, the emulator construction step that replaces $W$ with its low-rank counterpart $W^{\mathcal{E}}$ inevitably introduces a misalignment between these two models. As a result, the base model outputs seen by the LoRA modules are different during fine-tuning and inference. The difference is unexpected by the LoRA modules since they were optimized in the context of $W^{\mathcal{E}}$. To address this issue, we propose a novel LoRA correction algorithm and correct the misalignment explicitly, as shown in Fig. 3.
@@
where $\mathcal{V}_{\Lambda}$ denotes the subspace on which $\Lambda$ produces none-zero outputs—i.e., the region where it is active. It is worth noting that we enforce Eq. 3 only within $\mathcal{V}_{\Lambda}$, so that the correction has effects exclusively when the LoRA module is active. This design prevents unintended side effects in parts of the model unrelated to the LoRA module when doing correction, thereby preserving the integrity of the original model’s behavior.
Rewriting Eq. 3, we obtain an equivalent condition,
@@
Although this idea is conceptually straightforward, two practical challenges arise. First, the condition in Eq. 4 must hold for all inputs $x\in\mathcal{V}_{\Lambda}$. Second, there is no direct way to incorporate the correction term $\Delta_{x}$ into the LoRA weights $\Lambda=W_{A}W_{B}$.
The first challenge is resolved by leveraging two facts: the condition in Eq. 4 is linear, and $\mathcal{V}_{\Lambda}$ forms a subspace of the input feature space. As a result, it suffices to enforce the condition only on a basis of $\mathcal{V}_{\Lambda}$, denoted by $\beta_{\Lambda}$
@@
With the challenges addressed, we now formally describe the LoRA correction algorithm, as presented in Algorithm 1. We begin by preprocessing the original LoRA module $\Lambda=W_{A}W_{B}$ to obtain $W^{\prime}_{A}$ and $W^{\prime}_{B}$. Next, we select the columns of $W^{\prime}_{A}$ as the basis $\beta_{\Lambda}$ in Eq. 3. The correction term $\Delta$ is then computed by measuring the output difference between the original weight $W$ and the emulator weight $W^{\mathcal{E}}$ under this basis
@@
Each row of $\Delta$ represents the correction $\Delta_{a_{i}}$ corresponding to a basis vector $a_{i}$. Finally, we adjust $W^{\prime}_{B}$ to satisfy the condition in Eq. 5, compensating for the misalignment between $W$ and $W^{\mathcal{E}}$
@@
where the clamp function limits the norm of each correction vector $\Delta_{a_{i}}$ to be no more than a multiple $\lambda$ of the corresponding $b_{i}$. This hyperparameter $\lambda$ acts as a safeguard against excessive deviation, and we find it improves robustness in practice. The resulting corrected LoRA module, $\Lambda^{c}=W^{c}_{A}W^{c}_{B}$, is thus adapted to account for the misalignment introduced by training on the emulator and inferring on the original model.
Table 1: Performance comparisons of finetuning approaches on VQA with different memory costs. With InternVL2.5-8B as the original model, and 50% indicates finetuning on 4B models is conducted. Note that direct finetuning on InternVL2.5-8B (i.e., rows in gray) exceeds the memory budgets of all other finetuning methods and thus is not feasible under the same constraints.

### v2

@@ Abstract
Open-source foundation models have seen rapid adoption and development, enabling powerful general-purpose capabilities across diverse domains. However, fine-tuning large foundation models for domain-specific or personalized tasks remains prohibitively expensive for most users due to the significant memory overhead beyond that of inference. We introduce EMLoC, an Emulator-based Memory-efficient fine-tuning framework with LoRA Correction, which enables model fine-tuning within the same memory budget required for inference. EMLoC constructs a task-specific light-weight emulator using activation-aware singular value decomposition (SVD) on a small downstream calibration set. Fine-tuning then is performed on this lightweight emulator via LoRA. To tackle the misalignment between the original model and the compressed emulator, we propose a novel compensation algorithm to correct the fine-tuned LoRA module, which thus can be merged into the original model for inference. EMLoC supports flexible compression ratios and standard training pipelines, making it adaptable to a wide range of applications. Extensive experiments demonstrate that EMLoC outperforms other baselines across multiple datasets and modalities. Moreover, without quantization, EMLoC enables fine-tuning of a 38B model, which originally required 95GB of memory, on a single 24GB consumer GPU—bringing efficient and practical model adaptation to individual users. Project Page: hsi-che-lin.github.io/EMLoC
@@ 1 Introduction
General-purpose foundation models have demonstrated impressive zero-shot capabilities across a wide range of benchmarks [grattafiori2024llama, deepseekai2024deepseekv3technicalreport, Qwen2.5-VL, team2024gemma, chen2024expanding]. For real-world deployment, such as domain-specific tasks [lu2024multimodal, yue2023disclawllm] or personalized user behavior [yollava, shao2023character] , further customization by fine-tuning is still required. However, fine-tuning typically incurs significantly more memory overhead than inference [rajbhandari2020zero]. Consequently, if users have a fixed amount of available computing resources, they will be forced to choose between two unfavorable options. First, they can use a small model that fits within their memory budget for fine-tuning as shown in Fig.˜1(a), but this sacrifices the emergent capabilities [Wei2022EmergentAO] of larger models and underutilizes hardware during inference. Alternatively, they can opt for a large model that fully utilizes resources during inference but exceeds memory limits for fine-tuning as shown in Fig.˜1(b), making user-specific adaptation infeasible and potentially limiting performance in specialized applications. This paper addresses a central research question: Is it possible to design a fine-tuning strategy such that users can fine-tune a model under the same memory budget as inference?
@@
The memory cost of fine-tuning can be broadly attributed to three components: optimizer states, intermediate activations, and the model parameters themselves, as marked in Fig.˜1 with different colors. Initial efforts to reduce the memory usage of fine-tuning concentrated on the first two components. The first component, optimizer states, stores auxiliary information such as momentum and variance in the Adam optimizer [DBLP:journals/corr/Adam] for each trainable parameter. This overhead can be mitigated using parameter-efficient fine-tuning (PEFT) methods [hu2022lora, liu2024dora, yin2024parameter], such as LoRA [hu2022lora], which reduces trainable parameters by only updating lightweight low-rank matrices. The second component, intermediate activations, refers to the values retained during the forward pass for use in backpropagation. Previous works reduced it by modifying model architectures [sung2022lst, gomez2017reversible, liao2023make] or gradient checkpointing [chen2016training]. However, since the components reduced by both strategies still presented during fine-tuning, reducing them alone, without addressing the memory usage of model parameters, cannot fully close the memory gap between fine-tuning and inference (Fig.˜1(b)).
@@
Few recent works have taken initial steps toward addressing the third component, the memory cost of model parameters. For example, Offsite-Tuning [xiao2023offsite] proposes to drop some of the intermediate layers entirely during fine-tuning and to update only the top and bottom layers, thereby saving some memory occupied by model parameters. LORAM [zhang2025train] introduces row-pruning as a method to reduce model parameters during fine-tuning, allowing the updating of the unpruned rows in the intermediate layers. Despite their promise, these methods typically rely on memory-intensive full-model training. This entails a strong and often impractical assumption that well-resourced institutions must be involved, making them unfeasible for individual users. In addition, they impose strict constraints on which subsets of weights can be updated, limiting their flexibility and general applicability across tasks and architectures. Moreover, they overlook the fact that different base models are used during fine-tuning and inference, introducing potential mismatches that degrade performance.
@@
Figure 1: The dilemma caused by additional memory overhead during fine-tuning. (a) Users opt for a smaller 8B model, sacrificing emergent capabilities and underutilizing available hardware. (b) Use of a larger 26B model requiring memory exceeding the hardware limit even with LoRA [hu2022lora] and gradient checkpointing [chen2016training] techniques. (c) Our EMLoC utilizes a smaller model during fine-tuning, allowing the same budget for both training and inference.
@@
In this paper, we propose EMLoC, an Emulator-based Memory-efficient fine-tuning framework with LoRA Correction. EMLoC completely eliminates the memory cost gap between inference and training and can be carried out solely by individual users. The central idea is to perform fine-tuning on a low-rank model which we call the emulator so that the memory taken up by model parameters can be reduced as depicted in Fig.˜1(c). To make this approach effective, we introduce an efficient, downstream-aware emulator construction procedure based on activation-aware singular value decomposition (SVD), using a small subset of downstream data for calibration. This yields a lightweight emulator tailored to the downstream data, without requiring costly full-model training. Fine-tuning is then conducted entirely on the emulator without any assumptions about the fine-tuning process—any standard training pipeline can be used. However, since the LoRA modules are fine-tuned on the emulator rather than the original model, naively merging the LoRA modules into the original model will suffer from the misalignment between the full model and the compressed emulator. To mitigate this issue, we further introduce a novel LoRA correction algorithm, which adjusts the parameters to compensate such discrepancy during inference.
@@ 2.1 Reducing Optimizer State
There are mainly two ways to reduce the memory required for optimizer states: by reducing the number of trainable parameters [hu2022lora, yin2024parameter, zaken2021bitfit, zhang2023adalora, mahabadi2021parameterefficient, he2015delving, liu2024dora], or by reducing auxiliary information required per parameter [shazeer2018adafactor, li2023memory, zhao2024galore]. A prominent example of the first approach is LoRA [hu2022lora], which introduces trainable low-rank adapters into frozen pre-trained models to achieve parameter-efficient fine-tuning. For the second approach, Adafactor [shazeer2018adafactor] eliminates the need to store full-rank second-moment estimators by using a factored approximation. While these methods are effective, a non-trivial number of trainable parameters are required to maintain acceptable performance. As a result, the optimizer states associated with those parameters continue to contribute a non-negligible portion of the overall memory footprint.
@@ 2.2 Reducing Intermediate Activation
Another way to fine-tune with less memory overhead is to reduce memory from intermediate activations. Two explored strategies are shortening the backpropagation path [sung2022lst, mercea2024time, yin2024parameter, zhang2020side] and recomputing activations during the backward pass [chen2016training, gomez2017reversible, liao2023make]. An example of the first strategy is Ladder Side-Tuning [sung2022lst], where trainable modules are placed outside the backbone to avoid backpropagation through the backbone. [liao2023make] carefully design the position and initialization of the adapter modules so that the pre-trained model becomes reversible, which enables recomputing intermediate activation from the output during backpropagation. However, these methods cannot completely eliminate activation memory due to the fundamental need for gradients, and they typically incur additional computational overhead.
@@ 2.3 Reducing Model Parameters
Beyond optimizing optimizer states and intermediate activations, very recently reducing the memory footprint of the model parameters during fine-tuning has been explored. It is worth noting that quantization-based methods [dettmers2023qlora, zhang2024quantized] are largely orthogonal to our focus: they do not reduce fine-tuning memory relative to inference, and can therefore be combined with our method. An approach to reducing parameter memory during fine-tuning involves using a pruned model for training while maintaining a full model for inference [xiao2023offsite, zhang2025train]. For example, LORAM [zhang2025train] assumes that the model publisher provides a row-pruned and continually pre-trained variant of the original model. Users fine-tune this reduced model and then transfer the learned LoRA modules back to the full model. However, their setting requires continual pre-training of the pruned model, making it inaccessible to individual users. Furthermore, only the unpruned weights can be fine-tuned, limiting flexibility in selecting which parameters to adapt. In contrast, our work introduces a pipeline that can be performed by users individually and allows flexible selection of tunable parameters.
@@ 3 Method
Our proposed method, EMLoC, is a memory-efficient fine-tuning framework that enables individual users to fine-tune a model with the same memory cost as inference. An overview of EMLoC is shown in Fig.˜2. EMLoC consists of three main stages. First, we construct a downstream-aware lightweight emulator by performing activation-aware singular value decomposition (SVD) [wang2025svdllm] on a small calibration set drawn from the downstream data. Second, we fine-tune the emulator with LoRA modules using any standard training pipeline. The reduced parameter size of the emulator is crucial for enabling memory-efficient fine-tuning. Finally, to address the misalignment issue between the original model and the emulator, we present a novel LoRA correction algorithm to correct the learned LoRA modules, and transfer them back to the original model before inference.
@@ 3.2 Downstream-aware Emulator Construction and Fine-tuning
Specifically, we construct the set of emulator weights $\mathbb{W}^{\mathcal{E}}$ by applying the activation-aware SVD method proposed in SVD-LLM [wang2025svdllm] to each weight matrix $W\in\mathbb{W}$
@@
Owing to the specific design choices in the emulator, we can apply any standard PEFT techniques [hu2022lora, yin2024parameter, zaken2021bitfit] to fine-tune the significantly smaller emulator, achieving memory-efficient fine-tuning. In particular, if we aim to fine-tune a weight matrix $W_{i}$ in the original model, we instead insert a LoRA module next to $W^{\mathcal{E}}_{i}$ to fine-tune the emulator as in conventional frameworks. This allows us to use any existing training pipeline without modification, making our method a plug-and-play replacement for conventional LoRA fine-tuning, while further reducing memory usage.
@@ 5. $W^{c}_{B}\leftarrow W^{\prime}_{B}-\text{clamp}(\Delta,\lambda)$
To fully utilize the capacity of the original model during inference, we aim to transfer the LoRA modules $\Lambda$ fine-tuned on $\mathcal{E}$ back to the original model $\mathcal{M}$. However, the emulator construction step that replaces $W$ with its low-rank counterpart $W^{\mathcal{E}}$ inevitably introduces a misalignment between these two models. As as result, the base model outputs seen by the LoRA modules are different during fine-tuning and inference. The difference is unexpected by the LoRA modules since they were optimized in the context of $W^{\mathcal{E}}$. To address this issue, we propose a novel LoRA correction algorithm and correct the misalignment explicitly, as shown in Fig.˜3.
@@
where $\mathcal{V}_{\Lambda}$ denotes the subspace on which $\Lambda$ produces none-zero outputs—i.e., the region where it is active. It is worth noting that we enforce Eq.˜3 only within $\mathcal{V}_{\Lambda}$, so that the correction has effects exclusively when the LoRA module is active. This design prevents unintended side effects in parts of the model unrelated to the LoRA module when doing correction, thereby preserving the integrity of the original model’s behavior.
Rewriting Eq.˜3, we obtain an equivalent condition,
@@
Although this idea is conceptually straightforward, two practical challenges arise. First, the condition in Eq.˜4 must hold for all inputs $x\in\mathcal{V}_{\Lambda}$. Second, there is no direct way to incorporate the correction term $\Delta_{x}$ into the LoRA weights $\Lambda=W_{A}W_{B}$.
The first challenge is resolved by leveraging two facts: the condition in Eq.˜4 is linear, and $\mathcal{V}_{\Lambda}$ forms a subspace of the input feature space. As a result, it suffices to enforce the condition only on a basis of $\mathcal{V}_{\Lambda}$, denoted by $\beta_{\Lambda}$
@@
With the challenges addressed, we now formally describe the LoRA correction algorithm, as presented in Algorithm˜1. We begin by preprocessing the original LoRA module $\Lambda=W_{A}W_{B}$ to obtain $W^{\prime}_{A}$ and $W^{\prime}_{B}$. Next, we select the columns of $W^{\prime}_{A}$ as the basis $\beta_{\Lambda}$ in Eq.˜3. The correction term $\Delta$ is then computed by measuring the output difference between the original weight $W$ and the emulator weight $W^{\mathcal{E}}$ under this basis
@@
Each row of $\Delta$ represents the correction $\Delta_{a_{i}}$ corresponding to a basis vector $a_{i}$. Finally, we adjust $W^{\prime}_{B}$ to satisfy the condition in Eq.˜5, compensating for the misalignment between $W$ and $W^{\mathcal{E}}$
@@
where the clamp function limits the norm of each correction vector $\Delta_{a_{i}}$ to be no more than a multiple $\lambda$ of the corresponding $b_{i}$. This hyperparameter $\lambda$ acts as a safeguard against excessive deviation, and we find it improves robustness in practice. After applying Algorithm˜1 to each LoRA module independently, the resulting corrected weights, $\Lambda^{c}=W^{c}_{A}W^{c}_{B}$, is thus adapted to account for the misalignment introduced by training on the emulator and inferring on the original model.
4 Experiment
Table 1: Performance comparisons of finetuning approaches on VQA with different memory costs. With InternVL2.5-8B as the original model, and 50% indicates finetuning is conducted on 4B models. Note that direct finetuning on InternVL2.5-8B (i.e., rows in gray) exceeds the memory budgets of all other finetuning methods and thus is not feasible under the same constraints.